# OmniSAM: Omnidirectional Segment Anything Model for UDA in Panoramic Semantic Segmentation

### v2

@@ Abstract
Segment Anything Model 2 (SAM2) has emerged as a strong base model in various pinhole imaging segmentation tasks. However, when applying it to $360^{\circ}$ domain, the significant field-of-view (FoV) gap between pinhole ($70^{\circ}\times 70^{\circ}$) and panoramic images ($180^{\circ}\times 360^{\circ}$) poses unique challenges. Two major concerns for this application includes 1) inevitable distortion and object deformation brought by the large FoV disparity between domains; 2) the lack of pixel-level semantic understanding that the original SAM2 cannot provide. To address these issues, we propose a novel OmniSAM framework, which makes the first attempt to apply SAM2 for panoramic semantic segmentation. Specifically, to bridge the first gap, OmniSAM first divides the panorama into sequences of patches. These patches are then treated as image sequences in similar manners as in video segmentation tasks. We then leverage the SAM2’s memory mechanism to extract cross-patch correspondences that embeds the cross-FoV dependencies, improving feature continuity and the prediction consistency along mask boundaries. For the second gap, OmniSAM fine-tunes the pretrained image encoder and reutilize the mask decoder for semantic prediction. An FoV-based prototypical adaptation module with dynamic pseudo label update mechanism is also introduced to facilitate the alignment of memory and backbone features, thereby improving model generalization ability across different sizes of source models. Extensive experimental results demonstrate that our method outperforms the state-of-the-art methods by large margins, e.g., 79.06% (10.22$\uparrow$) on SPin8-to-SPan8, 62.46% (6.58$\uparrow$) on CS13-to-DP13.
@@ 1 Introduction
We conduct extensive experiments on the proposed OmniSAM framework and compare it with state-of-the-art (SoTA) methods [38, 42, 45] in Pinhole-to-Panoramic scenario. The results shows that our framework demonstrates superior performance over the SoTA methods by $\textbf{10.22}\%$ in indoor scenes and $\textbf{6.58}\%$ in outdoor scenes with trainable parameters less than $26$MB. In summary, our contributions can be summarized as: 1) We propose the OmniSAM framework, making the first attempt to explore the potential of SAM2 for the pinhole-to-panoramic semantic segmentation. 2) We introduce a dynamic pseudo label update strategy to mitigate the negative effects brought by fake pseudo labels during adaptation stage. 3) We propose a novel FPA module to effectively align features across domains and enhance the model’s robustness against distortions in panoramic images.
@@ 2.1 Segment Anything Model 2
SAM2 [26] is the second generation of the SAM [15], designed for instance-level segmentation in both image and video tasks. Trained on $50.9$K videos with $642.6$K masklets, SAM2 features a powerful image encoder and memory-driven modules to capture cross-frame relations. Research on SAM focuses on improving instance-level segmentation [35] and applying it to downstream tasks such as medical imaging [22, 25] and remote sensing [24]. Despite its ability to generate high-quality instance-level masks, SAM2 still lacks semantic knowledge. To address this, adapters and LoRA layers are introduced for domain fine-tuning [5, 4, 12], enabling knowledge transfer to improve performance in specific domains [46, 24]. In multi-modal semantic segmentation, several recent studies [17, 32, 23] have leveraged the pre-trained SAM/SAM2 backbone to extract high-dimensional representations from different modalities. These approaches introduce novel modules to fuse these features for optimal cross-domain representation. However, generalization to other domains still remains challenging. In this paper, we apply SAM2 for UDA in panoramic semantic segmentation. We propose a OmniSAM framework, which adapts SAM2’s architecture and memory modules for enhanced knowledge transfer.
@@ 4.2 Source Model Domain Gaps
We quantify the Pin2Pan domain gaps by comparing performance between standard pinhole (SPin8/CS13) and panoramic (SPan8/DP13) settings in both indoor and outdoor scenarios. As shown in Table 1, OmniSAM models consistently outperform the baseline. In indoor scenes, the OmniSAM models demonstrate strong generalization capabilities without requiring adaptation techniques. Notably, OmniSAM-L surpasses Trans4PASS+-S by over $13\%$ on indoor panoramic images, achieving $73.77\%$, indicating superior adaptation to panoramias. The tiny, small, and large variants of OmniSAM achieves even better performance on the target domain than on the source domain. These results may seem counterintuitive, as the model’s performance in the target domain even surpasses that in the source domain. This is very likely the results of the low-quality ground truth annotations of small objects (e.g. chair, sofa, table) in the target domain test set. Nevertheless, the significant improvement on this benchmark and visualizations still demonstrate the superiority of our model. For outdoor Pin2Pan test case, the Pin2Pan gaps are more pronounced, with Trans4PASS+-S experiencing a significant decline of $-23.04\%$ in mIoU, highlighting its challenge in handling panoramic distortions. In contrast, the OmniSAM models consistently reduce this gap, with OmniSAM-B showing the smallest drop of $-20.51\%$ in mIoU. Despite the domain gap, OmniSAM achieves superior performance in the target domain, with mIoU improvements ranging from $3.59\%$ to $5.21\%$. In outdoor Syn2Real scenario, OmniSAM-B outperforms Trans4PASS+-S by $2.34\%$.
@@ 4.3 Experimental Results
First, we evaluate our proposed framework under the SPin8-to-SPan8 setting. The results for source model are presented in Fig. 5 and Table 2, demonstrating both the source model capabilities and the effectiveness of adaptation. Given the source model performance in Table 1, the small, base, large variants of OmniSAM outperform the baseline model, 360SFUDA++, even without requiring adaptation techniques. Notably, the large OmniSAM variant achieves the highest mIoU score of $76.85\%$, demonstrating superior generalization in the indoor Pin2Pan task. Subsequently, we test our FPA method on both OmniSAM with memory attention w/ and w/o MA. The results indicate significant improvements on each model after bridging the knowledge gap. Our framework outstrips 360SFUDA++ by a margin of $+0.26\%$ to $+10.22\%$ in mIoU. Notably, OmniSAM-L with memory modules scores the highest $79.06\%$, achieving the SoTA performance on the indoor Pin2Pan task.

### v3

@@ Abstract
[…]
@@ 1 Introduction
We conduct extensive experiments on the proposed OmniSAM framework and compare it with state-of-the-art (SoTA) methods [38, 42, 45] in Pinhole-to-Panoramic scenario. The results shows that our framework demonstrates superior performance over the SoTA methods by $\textbf{10.22}\%$ in indoor scenes and $\textbf{6.58}\%$ in outdoor scenes with trainable parameters less than 2626MB. In summary, our contributions can be summarized as: 1) We propose the OmniSAM framework, making the first attempt to explore the potential of SAM2 for the pinhole-to-panoramic semantic segmentation. 2) We introduce a dynamic pseudo label update strategy to mitigate the negative effects brought by fake pseudo labels during adaptation stage. 3) We propose a novel FPA module to effectively align features across domains and enhance the model’s robustness against distortions in panoramic images.
@@ 2.1 Segment Anything Model 2
SAM2 [26] is the second generation of the SAM [15], designed for instance-level segmentation in both image and video tasks. Trained on 50.950.9K videos with 642.6642.6K masklets, SAM2 features a powerful image encoder and memory-driven modules to capture cross-frame relations. Research on SAM focuses on improving instance-level segmentation [35] and applying it to downstream tasks such as medical imaging [22, 25] and remote sensing [24]. Despite its ability to generate high-quality instance-level masks, SAM2 still lacks semantic knowledge. To address this, adapters and LoRA layers are introduced for domain fine-tuning [5, 4, 12], enabling knowledge transfer to improve performance in specific domains [46, 24]. In multi-modal semantic segmentation, several recent studies [17, 32, 23] have leveraged the pre-trained SAM/SAM2 backbone to extract high-dimensional representations from different modalities. These approaches introduce novel modules to fuse these features for optimal cross-domain representation. However, generalization to other domains still remains challenging. In this paper, we apply SAM2 for UDA in panoramic semantic segmentation. We propose a OmniSAM framework, which adapts SAM2’s architecture and memory modules for enhanced knowledge transfer.
@@ 4.2 Source Model Domain Gaps
We quantify the Pin2Pan domain gaps by comparing performance between standard pinhole (SPin8/CS13) and panoramic (SPan8/DP13) settings in both indoor and outdoor scenarios. As shown in Table 1, OmniSAM models consistently outperform the baseline. In indoor scenes, the OmniSAM models demonstrate strong generalization capabilities without requiring adaptation techniques. Notably, OmniSAM-L surpasses Trans4PASS+-S by over $13\%$ on indoor panoramic images, achieving $73.77\%$, indicating superior adaptation to panoramias. The tiny, small, and large variants of OmniSAM achieves even better performance on the target domain than on the source domain. These results may seem counterintuitive, as the model’s performance in the target domain even surpasses that in the source domain. This is very likely the results of the low-quality ground truth annotations of small objects (e.g. chair, sofa, table) in the target domain test set. Nevertheless, the significant improvement on this benchmark and visualizations still demonstrate the superiority of our model. For outdoor Pin2Pan test case, the Pin2Pan gaps are more pronounced, with Trans4PASS+-S experiencing a significant decline of $23.04\%$ in mIoU, highlighting its challenge in handling panoramic distortions. In contrast, the OmniSAM models consistently reduce this gap, with OmniSAM-B showing the smallest drop of $20.51\%$ in mIoU. Despite the domain gap, OmniSAM achieves superior performance in the target domain, with mIoU improvements ranging from $3.59\%$ to $5.21\%$. In outdoor Syn2Real scenario, OmniSAM-B outperforms Trans4PASS+-S by $2.34\%$.
@@ 4.3 Experimental Results
First, we evaluate our proposed framework under the SPin8-to-SPan8 setting. The results for source model are presented in Fig. 5 and Table 2, demonstrating both the source model capabilities and the effectiveness of adaptation. Given the source model performance in Table 1, the small, base, large variants of OmniSAM outperform the baseline model, 360SFUDA++, even without requiring adaptation techniques. Notably, the large OmniSAM variant achieves the highest mIoU score of $76.85\%$, demonstrating superior generalization in the indoor Pin2Pan task. Subsequently, we test our FPA method on both OmniSAM with memory attention w/ and w/o MA. The results indicate significant improvements on each model after bridging the knowledge gap. Our framework outstrips 360SFUDA++ by a margin of $0.26\%$ to $10.22\%$ in mIoU. Notably, OmniSAM-L with memory modules scores the highest $79.06\%$, achieving the SoTA performance on the indoor Pin2Pan task.